OpenStudy (anonymous):

2√2(√2-√10)+√3(√3+√15)

OpenStudy (anonymous):

your first job is to multiply out using the distributive law \[2\sqrt{2}\times \sqrt{2}-2\sqrt{2}\times \sqrt{10}+\sqrt{3}\times \sqrt{3}+\sqrt{3}\times \sqrt{15}\]

OpenStudy (anonymous):

some of these are easy, for example \(\sqrt{2}\times \sqrt{2}=2\) and \(\sqrt{3}\times \sqrt{3}=3\)

OpenStudy (anonymous):

this gives \[4-2\sqrt{20}+3+\sqrt{45}=7-2\sqrt{20}+\sqrt{45}\]

OpenStudy (anonymous):

then \(20=4\times 5\) so \(\sqrt{20}=\sqrt{4}\sqrt{5}=2\sqrt{5}\) so you have \[7-4\sqrt{5}+\sqrt{45}\] now repeat the same process with \(\sqrt{45}\) and finally combine like terms

OpenStudy (anonymous):

@satellite73 7-2√5 +3 √5 ?

OpenStudy (anonymous):

\[7-\sqrt{5}\]

OpenStudy (anonymous):

that should be your answer

OpenStudy (anonymous):

satellite deserve all the credit